OpenStudy (anonymous):

what is the distance between points ( 4, -5) and (0, 5)?

OpenStudy (anonymous):

I hope this helps :)

jimthompson5910 (jim_thompson5910):

Use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((0-4)^2+( 5-(-5))^2) d = sqrt((0-4)^2+( 5+5)^2) d = sqrt((-4)^2+(10)^2) d = sqrt(16+100) I'll let you finish.